OpenStudy (help!!!!):

Fan, medal, and testimonial if you help me with this problem!! How many numbers between 250 and 700 can be formed by using 2, 3, 4, 5, 6 with no repeats? My answer 66?? I need help on the steps

OpenStudy (help!!!!):

@jim_thompson5910 @mathmate @mathscience4ever @mathstudent55

jimthompson5910 (jim_thompson5910):

Focusing on just 3, 4, 5, 6 for the first digit, we have 4 choices for slot 1 That leaves 5-1 = 4 digits left over for slot 2 and 5-2 = 3 digits for slot 3 4*4*3 = 16*3 = 48 different 3 digit numbers that start with 3,4,5, or 6 --------------------------- if we have 2 in the first slot, then we can't have 3 or 4 in the second slot. Only 5 or 6 can be picked. That leaves 3 digits left over 1*2*3 = 6 there are 6 different permutations that start with 2 and are larger than 250 they are... 253 254 256 263 264 265 ------------------------------------------------------- in total, there are 48+6 = 54 different numbers that are between 250 and 700

OpenStudy (help!!!!):

could you explain how you got 6 permutations?